### v1

# Estimating Cognitive Effort from Functional Near-Infrared Spectroscopy (fNIRS) Signals using Machine Learning

Abstract.
The estimation of cognitive effort could potentially help educators to modify material to enhance learning effectiveness and student engagement. Where cognitive load refers how much work the brain is doing while someone is learning or doing a task cognitive effort consider both load and behavioral performance. Cognitive effort can be captured by measuring oxygen flow and behavioral performance during a task. This study infers cognitive effort metrics using machine learning models based on oxygenated hemoglobin collected by using functional near-infrared spectroscopy from the prefrontal cortex during an educational gameplay. In our study, sixteen participants responded to sixteen questions in an in-house Unity-based educational game. The quiz was divided into two sessions, each session consisting of two task segments. We extracted temporal statistical and functional connectivity features from collected oxygenated hemoglobin and analyzed their correlation with quiz performance. We trained multiple machine learning models to predict quiz performance from oxygenated hemoglobin features and achieved accuracies ranging from 58%–67% accuracy. These predictions were used to calculate cognitive effort via relative neural involvement and efficiency, which consider both brain activation and behavioral performance. Although quiz score predictions achieved moderate accuracy, the derived relative neural efficiency and involvement values remained robust. Since both metrics are based on the relative positions of standardized brain activation and performance scores, even small misclassifications in predicted scores preserved the overall cognitive effort trends observed during gameplay.
Machine learning, cognitive effort, relative neural efficiency, relative neural involvement, performance score, functional Near-Infrared Spectroscopy (fNIRS), brain signal, hemodynamic response, educational games
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: XXXXX; XX; isbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Interactive learning environmentsccs: Applied computing Learning management systemsccs: Applied computing Computer-assisted instruction
Figure 1. The overview of the proposed system (a) fNIRS data collected during quiz gameplay; (b) signals pre-processed and $\Delta HbO$ extracted; (c) statistical and connectivity features derived; Temporal feature engineering applied (d) ML models predict quiz score; (e) predicted score and $\Delta HbO$ used to estimate cognitive effort via relative neural efficiency and involvement.
1. Introduction
Cognitive effort (CE) refers to the level of cognitive load which is an individual experiences while learning (Getchell and Shewokis, 2023; Koiler et al., 2022). Cognitive load refers to how much the brain is working during a task (Getchell and Shewokis, 2023; Shewokis et al., 2015; Hasan et al., 2023). In many cases, we see that cognitive load becomes high, but this increase alone does not clearly tell us whether the person is deeply involved in the task or struggling with it. High brain activity can occur in both situations (Getchell and Shewokis, 2023; Shewokis et al., 2015; Hasan et al., 2023).
This is where the idea of cognitive effort becomes useful. Unlike cognitive load, which only shows how much the brain is working, cognitive effort considers both brain activation and performance together (Getchell and Shewokis, 2023; Koiler et al., 2022). When both brain activity is high and performance is also good, we interpret that as involvement. But if the brain is working hard and performance is still poor, it likely reflects burden. In other words, CE helps us differentiate between productive effort and mental overload. Even when performance looks efficient, CE can reveal whether that efficiency required excessive mental strain. By combining brain load and performance, we get a clearer picture of whether the effort was manageable or overwhelming(Getchell and Shewokis, 2023; Koiler et al., 2022).
CE is associated with increased neural activity during problem-solving, decision-making, and learning (Getchell and Shewokis, 2023; Koiler et al., 2022). When students are learning new or complex tasks, measuring CE helps us understand how hard their brains work and how to improve the design of the educational content. Relative Neural Efficiency (RNE) and Relative Neural Involvement (RNI) are two key metrics that can be used to explain CE with behavioral performance  (Getchell and Shewokis, 2023; Koiler et al., 2022; Reddy et al., 2022). These metrics are derived from hemodynamic responses in the prefrontal cortex (PFC), via functional Near-Infrared Spectroscopy (fNIRS), alongside with behavioral performance. The PFC is responsible for executive functions such as decision-making, and problem-solving, which are critical in educational contexts. RNE indicates how efficiently an individual processes information and performs a task with minimal CE. Higher RNE indicates low effort and high performance, and lower RNE suggests high effort and low performance. On the other hand, RNI measures motivation and engagement. High RNI indicates the user is more engaged and motivated during a task (Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023; Koiler et al., 2022). A balanced RNE and RNI represents individual’s engagement and efficiency during a task. Thus, understanding RNE and RNI may help to optimize learning strategies, improve engagement, and enhance cognitive performance.
Traditional methods of evaluating student efficiency and involvement rely on test assessments, self-reports, or reaction times, which are often subjective and inconsistent (De La Cruz et al., 2022). A few recent studies used hemodynamic responses to understand learning efficiency along with subjective measures and cognitive load  (Pan et al., 2020; Jeun et al., 2022; Oku and Sato, 2021). However, CE prediction using fNIRS and machine learning models remains an open research challenge.
By applying machine learning models to estimate performance from fNIRS signals, we evaluate the possibility of estimating RNE and RNI, allowing the detection of efficient learning. Our work introduces a novel dataset, that captures both fNIRS hemodynamic responses and quiz performance score during an educational quiz game from sixteen participants. We predicted performance score using various AI models. Based on this and fNIRS data we observed cognitive effort by estimating relative neural efficiency (RNE) and involvement (RNI). Instead of optimizing only for prediction accuracy, this study emphasizes interpreting brain signal patterns to better understand cognitive effort. Our research questions (RQs) are as follows:
  1. $RQ_{1}$
    Can machine learning models accurately predict performance scores from fNIRS-based features during educational gameplay?
  2. $RQ_{2}$
    Can predicted performance scores from ML model reflect cognitive effort meaningfully through derived RNE and RNI metrics?
The paper is organized as follows. section 2 reviews relevant literature. section 3 outlines our experimental methodology, detailing participants, apparatus, study procedure, cognitive effort measurement methods, data acquisition procedures, pre-processing techniques, and feature extraction frame work.  section 4 presents results. section 5 discusses findings about the research questions, addressing practical implications, limitations, and possible improvements. Finally, section 6 summarizes the study’s findings and highlights its educational contributions.
2. Related Works
fNIRS and Machine Learning in Brain Signal Analysis:
Machine learning models have been widely used on fNIRS-derived hemodynamic responses to support brain computer interface (BCI) tasks such as speech decoding (Cooney et al., 2022), motor imagery (Chiarelli et al., 2018; Ma et al., 2021), and finger movement classification (Khan et al., 2021; Ortega and Faisal, 2021; Jabbar et al., 2020). These models also help classify pain (Fernandez Rojas et al., 2024; Eken et al., 2024) and neurological disorders like Alzheimer’s (Cicalese et al., 2020; Kim et al., 2023), Parkinson’s (Quattrocelli et al., 2024; Guevara et al., 2024), ADHD (Shin et al., 2023; Zhao et al., 2021), and depression (Li et al., 2025). Researchers have used fNIRS with ML to detect fatigue (Tanveer et al., 2019), estimate cognitive load (Zhang et al., 2024), and classify cognitive tasks such as mental arithmetic (Wickramaratne and Mahmud, 2021; Zhang et al., 2023). CNN and LSTM models, in particular, have achieved high accuracy. For example, Wickramaratne and Mahmud reported 87.14% accuracy in classifying arithmetic tasks using CNN (Wickramaratne and Mahmud, 2021). Some studies combined EEG with fNIRS to improve BCI performance (Cooney et al., 2022; Ortega and Faisal, 2021; Khan et al., 2021).
fNIRS for Cognitive Workload, Fatigue, and Performance Prediction:
Beyond application in BCI and health sectors, researchers have increasingly explored how fNIRS-based models can predict cognitive workload and performance in real-world settings. fNIRS has been increasingly applied to predict cognitive workload and performance. Zhang et al. showed that DeepNet models can predict pilot workload in flight simulations using CNN-attention mechanisms(Zhang et al., 2024). Similarly, Grimaldi et al. proposed a CNN-LSTM model to predict perceptual load (Grimaldi et al., 2024a) and cognitive load (Grimaldi et al., 2024b) forecasting in piloting tasks. They stated that an LSTM-based approach demonstrated superior performance in short-term workload prediction (Grimaldi et al., 2024b). In another study, Gado et al. classified cognitive workload in NASA task-load simulations using logistic regression (Gado et al., 2023). Fatigue has also been investigated using fNIRS signals (Ma et al., 2025; Gado et al., 2023; Haroon et al., 2024; Varandas et al., 2022). In addition to workload and fatigue assessment, recent studies have explored the use of ML/DL for performance efficiency estimation, particularly in educational and learning environments. The following section discusses related works in this domain.
fNIRS in Performance and Efficiency Prediction:
While most studies focus on classifying cognitive workload, researchers have also begun investigating how ML/DL can predict performance efficiency and engagement based on fNIRS signals by monitoring neural activity in different learning and task-based environments. Jeun et al. used logistic regression to predict task difficulty based on PFC activity before training. TAfter three weeks of training, participants showed improved executive function and reduced PFC activation, indicating increased neural efficiency (Jeun et al., 2022).
Pan et al. applied ML techniques to classify instructional approaches based on brain-to-brain coupling between instructors and learners. They found that machine learning models could distinguish between scaffolding-based learning and explanation-based learning better than single-brain models that highlights the role of interactive engagement in learning efficiency (Pan et al., 2020). Another approach was proposed by Saikia, where k-means clustering has been used to classify individuals using fNIRS data during a working memory task (Saikia, 2023). However, the model faced challenges in generalizability due to individual differences in brain responses. Oku and Sato used fNIRS data to model with random forests and logistic regression to classify student engagement and predict quiz performance in online learning platforms (Oku and Sato, 2021).
Figure 2. (a) Experimental setup: Participant playing an educational game wearing fNIRS headband. This setup captures hemodynamic responses in the prefrontal cortex, with data acquisition managed through COBI software on a secondary laptop. (b) 16 channels used for collecting $\Delta HbO$
[…]
Figure 3: (a) Experimental setup: Participant playing an educational game on one laptop wearing fNIRS headband. The fNIRS device is attached to another computer, which captures hemodynamic responses in the prefrontal cortex, with data acquisition managed through COBI software on a secondary laptop. (b) We can see an image of a brain. and pointing 16 channel of pre-frontal cortex that we used for collecting $\Delta HbO$
3. Experiment
In this section, we briefly described the data acquisition process with the user study. Later, we explained the pre-processing steps and the features we extracted.
3.1. User Study
Figure 3 represents the steps of the user study. The participants used a unity-based educational game. The game was followed Sharmin et al. provided questions (Sharmin et al., 2024b) to evaluate participants’ understanding of the basic structure of graph concepts, such as nodes, edges, loops. Throughout the study, we recorded the participants’ quiz scores and hemodynamic responses from the fNIRS. The Qualtrics survey platform was used to collect responses to demographic questionnaire and pre and post-test questions. The study was approved by the institutional review board.
3.1.1. Apparatus
Figure 2 presents study setup and fNIRS headband orientation. We used two computers and an fNIRS device (Figure 2 (a)). We used an 18-channel headband fNIRS imager 2000S (fNIRS Device LLC, Potomac, MD, USA), 4 light emitters, wavelength ($\lambda$) 730-850 nm with average power $<1$ mW) sampled at 10 Hz. The detectors were separated by 2.5 cm with a penetrating depth of approximately 1.2 cm. This headband was placed on participants’ foreheads, and fNIRS collected data for 16 channels (Figure 2(b)). Channel 1-4 and channel 13-16 collectively called Lateral PFC (LPFC) and Channel 5-12 called Ventromedial PFC (VMPFC). Participants interacted with the educational quiz on a Dell XPS touchscreen laptop (L1) [Intel (R) Core(TM) i7-7700HQ processor at 2.80 GHz]. Another desktop computer (L2) [Intel(R) Core(TM) i7-10700T processor at 2.00 GHz] was connected to the fNIRS device. The laptop was used to collect fNIRS data using Cognitive Optical Brain Imaging (COBI) Studio Software. We also used fNIRSoft Software (Version 4.9) to apply filters on $\Delta HbO$ data and generate CSV file for further processing.
3.1.2. Study Procedure
Figure 3. Overview of the study procedures: Participants began with consent, demographic surveys, and a demo. A 20-second resting baseline preceded each 140-second task segment (4 questions). Rest periods were added between segments and sessions. Pre- and post-tests were collected using Qualtrics.
[…]
Figure 2: Overview of the study procedure. Overview of the experimental protocol. Participants first received an introduction to the study and completed a demographic questionnaire. A training session was provided to familiarize them with the experimental interface and task format. The main session began with a 20-second resting baseline, followed by the first task segment consisting of four questions. After a short 20-second rest, the second segment commenced. This was followed by a longer break and administration of post-task questionnaires. The same procedure was repeated for a second session after 10-12 minutes break.The entire protocol was designed to assess participant responses across structured task blocks interspersed with rest periods.
After signing the consent form and filling out the demographic questionnaires, the participants first experienced a demo version where the study was clearly explained and instructions about sitting still as much as possible (see  Figure 3).
After completing a ten-question pre-test, they watched the recorded video tutorial where they learned the basic terminologies of the graph. In our study, we divided our 16 questions into two sessions, with each session having two segments and each segment having four questions. Participants had up to 30 seconds to answer each question, and after each question, they had five seconds of feedback. So, each segment was 140 seconds $((4\times 30=120+4\times 5=20)=140)$ long. In between the two segments, there was a 20-second rest period. After the first session, they had a 6-10 minute break. After each session, the participants completed a post-test. We randomized the question order to minimize learning effects and control for difficulty bias across participants.
3.1.3. Cognitive Effort Measure
In this study, we aimed to estimate cognitive effort by observing the RNE and RNI in two quiz sessions. RNE represents the relationship between brain activity and task performance. While RNE measures neural efficiency, RNI measures engagement,how much voluntary effort is exerted. As each session has two segments, we have total four segments with four questions each.
To quantify cognitive effort, we first calculated the average of $\Delta HbO$ signals, representing oxygenated hemoglobin concentration, within the PFC for each learning segment. We then computed the overall group mean (GM) and standard deviation (SD) of both $\Delta HbO$ and behavioral performance scores to derive standardized z-scores (Getchell and Shewokis, 2023).
Standardizing both neural and behavioral measures serves two purposes: (1) it reduces individual variability, and (2) it brings physiological effort and performance onto a common scale, enabling meaningful comparison. This z-scoring process forms the basis of RNE and RNI, which are grounded in the theoretical framework of Paas et al. (Paas et al., 2003) and extended to neuroimaging by Shewokis and Getchell (Getchell and Shewokis, 2023; Shewokis et al., 2015; Koiler et al., 2022).
Specifically, $P_{z}$ denotes the standardized performance score, and $M_{z}$ represents the standardized inverse of mean $\Delta HbO$, based on the rationale that better learning should ideally occur with less cognitive effort. The inverse transformation of $\Delta HbO$ reflects this assumption, aligning with prior literature that interprets lower prefrontal activation during correct performance as greater neural efficiency.
To ensure numerical stability, especially in cases where all participants performed equally and the standard deviation of the score approached zero, we added a small constant $\epsilon=0.001$ to the denominator of the z-score formulas.
The final RNE and RNI metrics are computed using a Cartesian transformation:
(1) $\hskip 56.9055ptP_{z}=\dfrac{Score_{i}-Score_{\text{(GM)}}}{Score_{\text{(SD)}}+\epsilon}\hskip 28.45274pt$
[…]
(3) $\hskip 56.9055ptRNE=\dfrac{P_{z}-CE_{z}}{\sqrt{2}}\hskip 28.45274pt$
(4) $\hskip 56.9055ptRNI=\dfrac{P_{z}+CE_{z}}{\sqrt{2}}\hskip 28.45274pt$
This representation allows each participant’s effort profile to be plotted in a two-dimensional space, where RNE captures efficiency (high performance, low effort), and RNI reflects involvement (combined mental effort and engagement). This approach improves the interpretability of neural data in relation to cognitive learning processes.
3.2. Data Acquisition
We collected fNIRS data using COBI, and the fNIRS recording rate was 10 samples per second. The data was saved in fNIRS format. Each question was presented for 30 seconds, so each file contained 300 rows for each question corresponding to 16 optode ($\Delta HbO$) values. Sixteen participants answered 16 questions each, leading to 256 responses in the dataset ($16\times 16$). The dataset has 168 correct responses (class 1) and 88 incorrect responses (class 0). We used only $\Delta HbO$ because previous studies showed that $\Delta HbO$ signals are more sensitive to changes in blood flow in the participant’ cortex relative to $\Delta HbR$ signals (Gui et al., 2025; Liu et al., 2019; Pan et al., 2018). Since participants often answered within 20 seconds, only the first 200 rows per question were retained. The dataset contains 819,200 data points:
16 participants×16 questions×200 $16\text{ participants}\times 16\text{ questions}\times 200\text{ fNIRS}_{\text{points}}\times 16\text{ optodes}$.
3.3. fNIRS Signal processing using fNIRSoft
After data acquisition, we applied following steps on our fNIRS signals using fNIRSoft software (Biopac Systems Inc., [n. d.]). We checked raw light intensities and individual optodes visually. First, we rejected some noisy channels because of the failure to collect hemodynamic activity. It happened due to improper headband placement and contact between the sensors and the forehead. Next, we used a finite impulse response filter (20th order, Hamming window), a low pass filter, which removed the raw light intensity data at 0.1 Hz input from physiological signals, such as respiration and heartbeat. Then we applied the Modified-Beer Lambert rule, where fNIRS transforms light intensity measurements into changes in the concentrations of $\Delta HbO$ (Cakir et al., 2015; Çakır et al., 2016). According to the rule, a light’s attenuation depends on the medium’s characteristics it passes through (Cakir et al., 2015; Çakır et al., 2016). A portion of the near-infrared light used in fNIRS is absorbed by hemoglobin in the blood arteries in the brain. oxygenated hemoglobin ($\Delta HbO$) concentrations are associated with changes in brain activity. fNIRS calculates the changes in oxygenated hemoglobin, $\Delta HbO$, by quantifying how much light is absorbed (Cakir et al., 2015; Çakır et al., 2016). Then, the detrending filter was applied to data characterizing changes in concentration to remove drift in the data using linear parameters that convert the baseline slope to zero. We got oxygenated hemoglobin $\Delta HbO$ from the available biomarkers sent by Python script (Sharmin and Abrar, 2025; Sharmin et al., 2024a). At the end, we labeled the data using fNIRSoft software and saved them in CSV file format. Later, we extracted the quiz responses based on filenames. Then, we handled missing values by imputing with the mean per question and cleaned column names to prevent errors.
3.4. Feature Extraction Framework
The dataset contained fNIRS-derived $\Delta HbO$ signals from 16 optodes, including: Physiological data: optical density measurements (Optode 1-16); Experimental metadata: participant ID, question ID, question order, session; Labels: binary correct/incorrect answers. First, we identified optode-specific columns (Optode 1–16). For each optode, we computed Statistical Features (ST) and Functional Connectivity (FC).
Statistical Features (ST): A total of $128$ statistical features were extracted from $\Delta HbO$ signals, computed as $16$ optodes $\times$ $8$ statistical measures (mean, standard deviation, maximum, minimum, gradient mean, squared gradient mean, skewness, and kurtosis).
Functional Connectivity (FC) Features: For each question, we computed a $16\times 16$ Pearson correlation matrix using the 200-timepoint $\Delta HbO$ signals from all 16 optodes. These pairwise correlations capture functional connectivity between brain regions. To avoid redundancy, we extracted the 120 unique values from the upper triangle of the matrix (excluding the diagonal) as FC features.
Temporal Feature Engineering: Next, to capture temporal dynamics across the task, we computed delta ($\Delta$) features by subtracting the features of question $Q_{n-1}$ from $Q_{n}$ for each participant. This was done separately for both ST and FC features, resulting in $\Delta$ST and $\Delta$FC features. These deltas reflect changes in local brain activity and inter-regional connectivity between successive questions.
First, Data was sorted by Participant and Question_order. For each participant, temporal changes, $\Delta$-features for ST and FC were computed as:
(5) $\Delta feature=FeatureQ_{n}-Feature_{Q_{n}-1}$
To understand which brain regions most contribute to performance prediction, we analyzed temporal changes ($\Delta$ features) in ST and FC features across PFC subregions. FC captures how different brain regions interact over time. Changes in FC reflect network-level reconfiguration, which is critical for performance after break or task switch.
3.5. Model Implementation and Training Details
All models were implemented using scikit-learn (v1.4) and XGBoost (v2.0) in Python. Before training the models, we applied standardization (Z-score normalization) to ensure uniform feature scaling Finally, we used 5-fold cross-validation by applying GroupKFold $(n\_splits=5)$ to ensured participant-independent splits. Each fold consisted of a participant-wise split to ensure no overlap between train and test participants. Specifically, in each fold, approximately 80% of participants (13 or 12 out of 16) were used for training and the remaining 20% (3 or 4) for testing. This setup ensured participant-independent evaluation and avoided data leakage across folds. For preprocessing, all features were standardized using z-scores via StandardScaler.
The models and their key hyperparameters were as follows:
  • Logistic Regression: max_iter=1000
  • SVM (RBF): kernel=’rbf’, probability=True
  • KNN: n_neighbors=5
  • LDA: default parameters
  • Decision Tree: random_state=42
  • Random Forest: n_estimators=100
  • XGBoost: use_label_encoder=False, eval_metric=’logloss’
All models were evaluated using accuracy, precision, recall, and F1 score averaged across the 5 folds. Precision measures the proportion of correctly predicted positive observations to the total predicted positive observations. Recall evaluates the proportion of correctly predicted positive observations to all actual positive observations. A high recall indicates the model successfully captures most of the positive instances. F1-score is the harmonic mean of precision and recall. It is particularly useful when the class distribution is imbalanced.
To evaluate whether ML can reliably predict performance from fNIRS-derived features, we assessed multiple classifiers (e.g., Logistic Regression, SVM, Random Forest) across five types of features (Basic, statistical features (ST), Functional Connectivity (FC), ST+FC, and Temporal). These features are then used as input for various classification algorithms. Five different classifiers are used independently within a 5-fold stratified cross-validation framework: Logistic Regression (LR), SVM (RBF Kernel), K-Nearest Neighbors (KNN, k=5), Linear discriminant analysis (LDA), Decision Tree (DT), Random Forest (RF), XGBoost. The output of the classification is a binary prediction of correct (1) or incorrect (0) response.
4. Results
Table 1. Performance Comparison of Machine Learning Models. Higher represents better. Accuracy ranges from 58% to 67%
Feature Metric LR SVM KNN LDA DT RF XGBoost
Basic Accuracy 0.65 0.65 0.57 0.65 0.53 0.63 0.58
Precision 0.53 0.46 0.53 0.53 0.51 0.58 0.54
Recall 0.65 0.65 0.57 0.65 0.53 0.63 0.58
F1 Score 0.56 0.52 0.54 0.57 0.52 0.59 0.54
ST Accuracy 0.55 0.65 0.58 0.51 0.52 0.62 0.62
Precision 0.55 0.43 0.55 0.53 0.54 0.58 0.61
Recall 0.55 0.65 0.58 0.51 0.52 0.62 0.62
F1 Score 0.53 0.52 0.54 0.52 0.52 0.57 0.60
FC Accuracy 0.55 0.62 0.58 0.53 0.48 0.61 0.58
Precision 0.57 0.45 0.54 0.58 0.50 0.61 0.55
Recall 0.55 0.62 0.58 0.53 0.48 0.61 0.58
F1 Score 0.55 0.51 0.55 0.54 0.49 0.56 0.56
ST+FC Accuracy 0.56 0.64 0.58 0.51 0.53 0.67 0.65
Precision 0.58 0.52 0.59 0.58 0.54 0.65 0.63
Recall 0.56 0.65 0.58 0.51 0.53 0.67 0.65
F1 Score 0.56 0.53 0.56 0.52 0.54 0.61 0.62
Temporal Accuracy 0.62 0.64 0.65 0.50 0.66 0.56 0.59
Precision 0.63 0.68 0.63 0.57 0.68 0.59 0.57
Recall 0.62 0.64 0.65 0.50 0.66 0.56 0.59
F1 Score 0.62 0.55 0.56 0.52 0.61 0.56 0.54
4.1. Model Evaluation
From  Table 1, we can see that while the models achieved prediction accuracies ranging from 58% to 67%, which is above the random baseline, the performance suggests room for improvement in model robustness and feature refinement. For instance, the Random Forest classifier achieved an accuracy of 67% with ST+FC features, while SVM shows consistent performance across (62%-65%) feature types. Recall remained consistently moderate-to-high (up to 0.68) and indicates that the models were able to correctly identify a substantial proportion of correct and incorrect answer. These findings collectively demonstrate that ML models can feasibly predict task performance from brain signal-derived features, without relying on any specific model being best. No statistically significant differences were found between classifiers ($Wilcoxon,p>0.05$). This supports our focus on interpretability over performance optimization. The consistency across models suggests that fNIRS features encode robust cognitive signals that can be captured similarly by different classifiers. We also applied deep learning models such as CNN (60.05%), GRU (63.75%), LSTM (64.68%), and BiLSTM (63.43%) models. However, due to the limited size of our dataset—only 256 samples (16 participants × 16 questions)—the models tended to overfit, and validation accuracy did not significantly improve over traditional ML methods. The relatively high parameter count of DL architectures, combined with the small dataset, hindered generalization, which aligns with past findings on DL sensitivity to data volume in neuroimaging studies.
4.2. Estimation of Cognitive Effort
In this study, our focus was to find if predicted score using AI, can be helpful to estimate cognitive effort. To represent cognitive effort we have calculated relative neural efficiency (RNE) and neural involvement (RNI) using equation 4. These metrics provided a deeper understanding of participants’ cognitive states across task segments. We selected one test case that included three participants: P8, P11, and P16. we quantified similarity using Mean Absolute Error (MAE) and Pearson correlation actual and predicted RNE and RNI values across task segments for three test participants. The very low MAE (0.29 for RNE and 0,38 for RNI) and near-perfect correlation $(r=0.999)$ suggest strong alignment that the model’s predicted RNE and RNI values closely follow the actual measurements.
To understand participants’ efficiency and involvement we plotted Z-score standardized $\Delta HbO(M_{Z})$ in X axis and Z-score standardized performance score $P_{z}$ in Y axis. Figure 4 shows participant-wise RNE and RNI in a Cartesian plot. The diagonal from bottom-left to top-right marks the efficiency line—points above indicate high neural efficiency. The opposite diagonal represents involvement—points above it indicate high neural involvement. This divides the space into four cognitive states: HE+HI (High Efficiency, High Involvement), HE+LI (High Efficiency, Low Involvement), LE+HI (Low Efficiency, High Involvement), and LE+LI (Low Efficiency, Low Involvement). Predicted coordinates are plotted alongside actual values for participants P8, P11, and P16 across four task segments.
Out of the 12 predicted coordinates, 9 matched the actual cognitive states when visualized in the RNE-RNI Cartesian plot. Although the accuracies of the machine learning models are moderate, the plotting suggests strong interpretability in terms of cognitive state estimation. In our framework, RNE and RNI are evaluated not only by their values but also by their relative position and distance from the efficiency and involvement diagonals. Therefore, even when the exact predicted RNE/RNI values deviate from ground truth, the quadrant-wise agreement in most cases suggests that the model captures broad cognitive states rather than precise individual metrics. In segment 1, all participants showed low neural efficiency but varied involvement which often associated with the early learning phase. Segment 2 showed clustering near the origin for most participants, implying a moderate-effort transitional state. Segment 3 revealed decreased effort for P11 and P16, possibly prompting the subsequent long break. P8, who reported a headache after segment 2, exhibited higher cognitive effort in segment 3. Finally, in segment 4, cognitive effort generally declined again after a short rest, with participants transitioning toward lower-effort zones.
Figure 4. Cartesian representation of predicted vs. actual cognitive states. X-axis: standardized $\Delta HbO$ ($M_{Z}$); Y-axis: standardized performance score ($P_{Z}$). Circles = actual scores; crosses = predicted scores. Quadrants represent cognitive states: HE+HI (High Efficiency, High Involvement), HE+LI, LE+HI, LE+LI.
[…]
Figure : Predicted and actual performance scores across four task segments for each participant. Circles represent actual scores, and crosses represent predicted scores. The X-axis indicates Cognitive Effort (CE), and the Y-axis shows the Performance Score. Each quadrant corresponds to a cognitive state: High Efficiency+ High Involvement (HE+HI), High Efficiency+ Low Involvement (HE+LI), Low Efficiency+ High Involvement (LE+HI), and Low Efficiency+ Low Involvement (LE+LI).
These results highlight that our model’s predictions are cognitively meaningful, even when numeric prediction errors exist, supporting its utility for estimating learner state in real-time applications.
5. Discussion
The aim of this study is to estimate cognitive effort during educational tasks by analyzing fNIRS-derived hemodynamic responses and predicted performance scores. Our findings highlight that model predictions, even with moderate accuracy, can reveal meaningful cognitive trends when analyzed through RNE and RNI. We used various machine learning models to predict quiz performance from brain signals and then we derived Relative Neural Efficiency (RNE) and Relative Neural Involvement (RNI) to interpret learner efficiency and involvement. The contribution of this work is that it shows that cognitive effort can be meaningfully measured using a combination of brain signals and machine learning-based performance prediction. These findings might influence future adaptive learning systems.
For $RQ_{1}$:, The results of this study demonstrate that task performance prediction using fNIRS-derived brain features is not only feasible but also effective across a variety of feature types and classifiers. While our focus was not on identifying the single best-performing model, the consistent performance across classifiers confirms that fNIRS signals—especially when processed into spatial, functional, or temporal representations—contain rich cognitive information. The strong model generalization across ST, FC, and delta (Temporal) features highlights that both static brain activation and dynamic changes during tasks contribute meaningfully to performance estimation. There was no statistically significant performance difference. This consistency suggests that the cognitive patterns embedded in fNIRS signals are inherently stable and interpretable, regardless of the specific model architecture.
For $RQ_{2}$, these derived measures help interpret learners’ cognitive effort, involvement, and efficiency during educational gameplay. Even with moderate accuracy, the predicted CE trends followed actual RNE/RNI patterns. Since we use aggregated and z-scored values for CE estimation, occasional false positives/negatives have a limited impact on the overall cognitive state trends. The derived RNE/RNI values smooth out individual prediction errors. Also, to observe cognitive effort patterns, we used a Cartesian plot. We observed the predicted values most of the time were aligned with the actual values in the same quadrant of the Cartesian plot. This spatial overlap indicates that the model was able to capture similar patterns of cognitive effort behavior. Therefore, the proposed model not only supports cognitive effort estimation but also aids in the interpretation of cognitive state trends across different learning segments.
However, future improvements can reduce this noise further. Our goal was not to maximize classification accuracy alone, but to enable interpretable CE estimation using RNE and RNI. We plan to explore more robust or adaptive machine and deep learning architectures in future work. Cognitive effort is highly individualized. While our model captures general trends, it does not yet account for individual baselines or variability in fNIRS responses. Adaptive or personalized calibration may improve robustness in future work.
6. Conclusion
This study demonstrates a method for estimating and interpreting cognitive effort through performance scores predicted by various machine learning models and oxygenated hemoglobin ($\Delta HbO$) in the brain collected using fNIRS. This work demonstrates that AI models can support interpretable cognitive effort estimation, going beyond performance to inform adaptive learning design. Our findings confirm that ML-predicted performance, combined with fNIRS data, provides a viable estimate of neural efficiency and involvement. We observed the positive correlation between the brain feature dynamics and performance score in an educational game. This predicted score was used to estimate cognitive effort by measuring relative neural efficiency and involvement. The machine learning models showed accuracy ranges from 58% to 67%. After estimating RNE and RNI, we observed that the mean absolute error is very low between actual and predicted RNE/RNI values, which suggests that predicted RNE and RNI closely follow the actual RNE and RNI. This approach paves the way for a personalized educational setup that can dynamically adapt to learners’ cognitive states, potentially adjusting to classroom and remote educational settings. By deriving RNE and RNI from predicted performance and effort, we show meaningful trends tied to cognitive states. The findings highlight the model’s potential in adaptive learning contexts and highlight future directions in real-time neuro-adaptive systems.
Acknowledgments
We wish to express our gratitude to the study participants and lab members.